OpenStudy (anonymous):

HELP! MEDAL :D City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was –3° C. What was the temperature of City A and City B on that day? City A was 5° C, and City B was 4° C. City A was 3° C, and City B was –1° C. City A was 8° C, and City B was –3° C. City A was 5° C, and City B was –5° C.

OpenStudy (anonymous):

@ganeshie8 @phi @wio

OpenStudy (anonymous):

@dan815 @TheSmartOne @kropot72 @sammixboo @StudyGurl14

OpenStudy (haleyelizabeth2017):

So first, we are going to need to come up with equations. The first one, 4 times the temperature of a is going to be written as 4a, okay? and then 8 degrees more than 3 times the temperature of city b is going to be written as 3b+8. Does this make sense to you?

OpenStudy (haleyelizabeth2017):

And since that equals 4a, we can finish the first equation by writing 4a=3b+8

OpenStudy (anonymous):

Alright

OpenStudy (haleyelizabeth2017):

What do you think the second equation is going to be?

OpenStudy (anonymous):

For?

OpenStudy (haleyelizabeth2017):

For this problem there will be a total of 2 equations. We found the first one and what I need you to do is try to find the other one for "The temperature of City A minus twice the temperature of City B was –3° C"

OpenStudy (haleyelizabeth2017):

_-__=-3 is how we will set it up

OpenStudy (anonymous):

A - 2b = -2

OpenStudy (anonymous):

?

OpenStudy (haleyelizabeth2017):

Close. a-2b=-3

OpenStudy (anonymous):

My Caps :(

OpenStudy (anonymous):

ohh yeah sorry my bad with the -2

OpenStudy (haleyelizabeth2017):

lol it's okay :) next, we need to isolate the a in the equation we just came up with by adding 2b to both sides, okay? can you do that really quick?

OpenStudy (anonymous):

forgot how to isolate.... could you help me? explain :)

OpenStudy (haleyelizabeth2017):

Okay so we have: a-2b=-3 +2b +2b ----> since we have -2b, we add it to both sides to cancel it on the left side. ----------- a=2b-3

OpenStudy (anonymous):

AHH! lol I get it now

OpenStudy (haleyelizabeth2017):

Okay :) great

OpenStudy (haleyelizabeth2017):

Now we plug in (2b-3) for a in the first equation, 4a=3b+8

OpenStudy (haleyelizabeth2017):

So 4(2b-3)=3b+8 Do you know how to distribute?

OpenStudy (anonymous):

4 into 2b and 4 into 3?

OpenStudy (haleyelizabeth2017):

yep :)

OpenStudy (anonymous):

wait.. you need to add or multiply?

OpenStudy (haleyelizabeth2017):

multiply

OpenStudy (anonymous):

8b - 12

OpenStudy (haleyelizabeth2017):

Ding ding ding! We have a winner! So now that we have 8b-12=3b+8 we need to subtract 3b from both sides to cancel it out on the right side. 8b-12=3b+8 -3b -3b ------------- 5b-12=8 next, we need to add 12 to both sides. Do you want to try it out this time?

OpenStudy (anonymous):

17b = 20?

OpenStudy (haleyelizabeth2017):

Nope. Close though. You can't add it to 5b because they are not like terms. And when you add them, it just cancels on the side you are removing it from

OpenStudy (haleyelizabeth2017):

So it just stays as 5b=20

OpenStudy (haleyelizabeth2017):

So what is 20/5?

OpenStudy (anonymous):

4

OpenStudy (haleyelizabeth2017):

Great! Now we have b=4 because we divided both sides by 5 to isolate the b :)

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

@haleyelizabeth2017 ? is it A

OpenStudy (haleyelizabeth2017):

Lets solve for a first then we will know :)

OpenStudy (anonymous):

Alright

OpenStudy (haleyelizabeth2017):

Okay, now that we have b, we can plug it in to the equation. We had already isolated a in the second one so lets just plug it into that one :) a=2b-3

OpenStudy (haleyelizabeth2017):

So we do this a=2(4)-3

OpenStudy (haleyelizabeth2017):

What is 2*4? Then what is that minus 3?

OpenStudy (anonymous):

5

OpenStudy (haleyelizabeth2017):

Ding, ding, ding, we have yet another winner!! :D good job! Now we have finished :) so yes indeed, it is A :D

OpenStudy (anonymous):

:D

OpenStudy (haleyelizabeth2017):

Do you need help with any more?